# Self-Calibrating Gaussian Splatting for Large Field of View Reconstruction

### v1

@@ 3.3 Cubemap for Large FOV
In order to apply our method to cameras with larger FOV, we extend the single planar perspective projection to a cubemap projection [61, 35]. Mathematically, single planar projection requires upsampling in the peripheral regions, and the sampling rate increases drastically as the FOV approaches 180. In contrast, rendering with a cubemap maintains relatively uniform pixel density from the image center to the edges, making it ideal for wide-angle rendering.
@@
Given the parameters estimated from SfM [58], the parametric model is then applied to undistort the raw image into perspective images, as shown in the blue box in Fig. 2(a). These undistorted images are then reconstructed through perspective-based pipelines like NeRF [48] or 3D Gaussian Splatting (3DGS) [37]. However, this process stretches the pixels in the peripheral regions, and the effect becomes more pronounced when undistorting into larger FOV perspectives, as shown in the yellow box. More specifically, the stretching rate of each pixel is defined by the inverse of Eq. 3, which exhibits a trend similar to , where is the FOV angle from the center in the raw image. When the FOV for the undistorted image is 110 (as in the blue example), the upsampling rate from the blue circle to the box is around 1.4. However, when the FOV increases to 170 like the yellow one, this rate rises to 11.4, which inevitably sacrifices a significant amount of high-frequency information for reconstruction.
Moreover, to preserve central details, the resolution of undistorted images needs to be higher, as the pixel density at the center of the raw image should ideally match that of the undistorted perspective. For example, when undistorting a 180 fisheye image in Fig. 2(a), the resulting perspective image in the yellow region would have an extremely high resolution, making it computationally infeasible. A common solution is to crop away the periphery, following COLMAP’s approach [58], but this strategy contradicts our intent of using a fisheye camera to capture wide-angle information.
@@
Inspired by the representation of environment maps using cubemaps in computer graphics [24], we propose representing extreme wide-angle renderings with cubemap projections, each covering 90 FOV and oriented orthogonally to one another, as illustrated in Fig. 2(b). By resampling across the cubemap faces, we can render perspective or distorted images with FOVs even larger than 180. Fast rasterizations are first applied to obtain each face of the cubemap. For each rendered pixel, we look up its corresponding position in the constructed cubemap. Our hybrid distortion field then resamples from the lookup table to achieve the distorted rendering. In practice, it is not necessary to render all faces of the cubemap at once as the choice of number of cubemaps may differ for different FOV camera lenses.
The resampling step only involves a simple coordinate transformation along with our hybrid field distortion. The distance from the shared camera center to each plane is normalized to 1. To render a pixel outside the 90 FOV at where and , for instance, the pixel on the right face can be obtained as in camera coordinates, looking toward the right side. When considering lens distortion, the sampling mapping is distorted according to Eq. 10, altering the lookup on the right face to , where . By doing so, the entire distortion field can be directly applied to the cubemap for large FOV rendering. The entire resampling process is fully differentiable, making it directly applicable in our hybrid distortion field as a plug-and-play module.
@@
Vanilla Gaussian Splatting builds an ordered set of Gaussians before alpha blending [37]. The Gaussians are sorted based on depth (i.e., the orthogongal projection distance to the image plane). This strategy is valid if a single Gaussian is not projected onto two different faces. However, with the cubemap, Gaussians can span the boundary between two faces, resulting in multiple projections, and the depth ordering is different between the two faces, leading to intensity discontinuities at the boundary. To resolve this, we replace the original sorting by depth with sorting by distance from the camera center to each of the Gaussians. This change ensures that the rasterization order is consistent across all faces of the cubemap, thereby alleviating intensity discontinuities. Due to the affine approximation used in Eq. 5, the 2D covariance of the Gaussians at the cubemap face boundary still has a slight influence on the final render, which we discuss in Sec. 5.
@@
Dense-Grid | 0.511 | 15.14 | 0.525 | 0.316 | 13.29 | 0.581 | 0.513 | 15.36 | 0.517 | 0.556 | 16.02 | 0.489 | 0.556 | 16.03 | 0.489 | 0.381 | 13.74 | 0.527 |
Sparse-Grid | 0.532 | 15.79 | 0.442 | 0.529 | 16.06 | 0.513 | 0.339 | 14.19 | 0.560 | 0.535 | 16.29 | 0.502 | 0.576 | 16.58 | 0.475 | 0.419 | 15.10 | 0.482 |
@@
4 Results
In this section, we evaluate our method on both synthetic and real datasets. We will start by describing the dataset we use and the process taken to curate these datasets, followed by comprehensive evaluations.
Synthetic Data.
We customized a camera module in the Mitsuba ray tracer [33] to incorporate camera parameters derived from professional DSLR lenses and profiles from the open-source Lensfun database [1]. For scene setup, we rendered three large indoor scenes and four object-centric scenes using 3D assets from [9]. We created a held-out test set where each scene was rendered with paired perspective and fisheye cameras to accurately evaluate the performance of undistortion. To test the efficiency of fisheye captures, we created a training set where the number of fisheye images (i.e., 100) is half that of the narrower FOV perspective images. We applied various levels of lens distortion to the object-centric scenes to evaluate our lens model’s adaptability across diverse camera setups. Additionally, we introduced synthetic distortions, including both radial and tangential components, to images from the LLFF dataset [47]. Further details are provided in the supplementary material.
Real-World Data.
We captured several real-world datasets using different uncalibrated cameras to test our method, consisting of casual walk-around video footage with a camera that has a field of view of approximately 150. Additionally, we evaluated our method on the existing benchmark FisheyeNeRF dataset [34], which consists of highly distorted images captured with ultra-wide-angle fisheye lenses featuring a field of view of around 120. The dataset includes a variety of indoor and outdoor scenes with complex lighting and geometric details.
Room1 | ![]() |
Room2 | ![]() |
(a) 3DGS w/ Small FOV | (b) Ours | (c) Fisheye-GS | (d) Ours | (e) GT |
Garden | ![]() |
Studio | ![]() |
(a) 3DGS w/ COLMAP | (b) Ours (Undistort) | (c) Fisheye-GS | (d) Ours (Fisheye) | (e) Fisheye GT |
Method | Mitsuba Kitchen | Mitsuba Room1 | Mitsuba Room2 | ||||||
SSIM | PSNR | LPIPS | SSIM | PSNR | LPIPS | SSIM | PSNR | LPIPS | |
3DGS [37] | 0.470 | 11.09 | 0.435 | 0.595 | 15.27 | 0.406 | 0.897 | 30.88 | 0.155 |
Ours | 0.794 | 27.56 | 0.272 | 0.686 | 26.27 | 0.314 | 0.895 | 33.21 | 0.172 |
Fisheye-GS [46] | 0.601 | 14.30 | 0.485 | 0.549 | 15.54 | 0.548 | 0.739 | 17.97 | 0.355 |
Ours | 0.886 | 30.72 | 0.146 | 0.842 | 28.46 | 0.180 | 0.929 | 31.16 | 0.095 |
4.1 Comparisons with Traditional Lens Models
We evaluate the effectiveness of our hybrid distortion field (Eq. 10) in modeling substantial fisheye distortions. We use the FisheyeNeRF dataset [34], which consists of full-frame fisheye captures from two indoor and four outdoor scenes. Instead of undistorting images during preprocessing, we directly optimize Gaussians from uncalibrated fisheye photos, simultaneously accounting for camera distortion while optimizing camera parameters and 3D scenes together. We compare our method against four baselines: 1) Vanilla Gaussian Splatting (3DGS); 2) a discrete grid-based method similar to Jeong et al. [34] in both dense (Grid-Dense) and sparse (Grid-Sparse) versions; 3) Fisheye-GS [46], which modifies the 3DGS projection model to a specific type of fisheye parametric camera model; and 4) optimization of the omnidirectional camera model from ADOP [55], which we have re-implemented on GS.
Our method consistently outperforms the baselines in all image metrics, as shown in Tabs. 2 and 1. The baseline methods struggle at the four corners of the images, which can potentially be attributed to using too few parameters to model a complex fisheye camera and bad initializations from SfM, as illustrated in Fig. 4, Fig. 5 (c) and (d), and Fig. 6 (c) and (d). Our method can effectively model lens distortion, resulting in improved reconstruction quality.
Type | Method | PSNR | SSIM | LPIPS |
Mitsuba Indoor | Vanilla-GS [37] | 19.08 | 0.654 | 0.332 |
Ours (Single Planar) | 24.10 | 0.676 | 0.312 | |
Ours (Cubemap) | 29.01 | 0.792 | 0.253 | |
Mitsuba Objects | Vanilla-GS [37] | 25.37 | 0.923 | 0.121 |
Ours | 28.00 | 0.932 | 0.093 |
4.2 Large FOV Reconstruction.
To verify that large FOV cameras provide better reconstruction coverage than narrower FOV perspective cameras, we evaluate our method on both synthetic Mitsuba scenes and real-world captured scenes.
Our method, directly reconstructed from uncalibrated fisheye images, delivers superior reconstruction quality despite using only half the number of images required by narrower FOV cameras, as the fisheye images capture a larger area in each shot. Quantitatively, we report reconstruction evaluations on the hold-out camera set, as shown in the first two rows of Tab. 2. Reconstruction directly from fisheye captures using our pipeline significantly outperforms the perspective-based reconstruction supported by vanilla 3DGS [37]. As visualized in Fig. 5a and b, our method achieves better coverage in Room1, where all cameras are oriented toward the sofa. Additionally, our method achieves comparable reconstruction with fewer artifacts in a walk-around setting like Room2, even with fewer images.
We further evaluate our method on real-world fisheye captures with over 150 FOV. As a baseline, we apply COLMAP to undistort images from our dataset. We then apply vanilla 3DGS using the undistorted images and our method directly on the raw images, with camera extrinsics and intrinsics estimated by COLMAP. The results are shown in Fig. 6. We find that COLMAP tends to excessively crop the peripheries far from the center of the image, resulting in a significant loss of information (first column). Our approach avoids laborious pre-calibration while retaining details of the original captures, including those in highly distorted areas where information is typically lost during pre-undistortion processes. By optimizing the hybrid distortion field with a multi-planar projection, we are able to utilize the full field of view of fisheye captures, as shown in Fig. 6 (a) and (b).
@@ 4.1 Comparisons with Traditional Lens Models
4.3 Adaptability to Different Lens Distortions.
@@
To verify that our method can model various types of lens distortion, we create synthetic distortions to the images in the LLFF dataset [47]. This simulation mimics various real-world distortions, including radial and tangential distortions, using camera parameters derived from professional DSLR lenses profiled in the open-source Lensfun database. By applying these parameters, we generate the combination of different distortion as shown in the Trex scene. Fig. 7 shows that our method can accurately model various types of camera distortion without the need for manual calibration or access to the physical camera.
4.4 Optimization of Camera Parameters.
Our pipeline also supports camera parameters, including intrinsic and extrinsic optimization, during the reconstruction. The optimization is implemented within native CUDA kernela, making it efficient. Since the main focus is to handle distortion, we leave more experiments and verification of this capability in the supplementary material.
4.5 Ablation
In this section we validate the necessity of using cubmap resampling for enabling wide FoV reconstruction. We compare the result obtained by our hybrid distortion field using single planar project and using cubemap resmapling. We evaluate these two configurations on three wide-angle indoor scenes (i.e. Mitsuba Indooor). As presented in Tab. 3, cubemap projections yield significant improvements for 180° fisheyes. Note that both versions of our method outperform the baseline, showing the advantage of using the hybrid distortion field. Additional visualizations are provided in the supplementary.
@@ 5 Discussion
This paper presents a method for optimizing Gaussian splat representations while self-calibrating camera parameters and lens distortion. Our approach enables the use of large field-of-view captures to achieve efficient and high-quality reconstruction without cumbersome pre-calibration.
@@
In this work, we do not address the phenomenon of entrance pupil shift present in most fisheye lenses: rays in the center of the field of view converge at an entrance pupil deeper in the lens than rays at the periphery, which differs from the type of lens distortion we are modeling. In contrast, entrance pupil shift, though insignificant when the scene is distant, can cause splat misalignments for near-field scenes, as the shift can easily be on the order of half a centimeter for full-frame lenses. Additionally, we did not model vignetting (radial brightness falloff) or chromatic aberration. Extending our methods to comprehensively model these effects is a promising direction for future work. Also, since 2D covariances depend on the look-at directions of the cubemap faces onto which the 3D Gaussians are projected, Gaussians located at the boundaries of adjacent cubemap faces will have different 2D covariances, resulting in slight intensity discontinuities. We observe that simply changing the previous depth sorting strategy to distance sorting can largely mitigate these discontinuities. A potentially more elegant solution for future work could be to project the covariances orthogonally to the viewing ray rather than onto the cubemap face.
@@ 1 Supplementary Video
We provide a video to better compare our methods with baselines. Our video is organized into three different parts.
The first part is a comparison between our method and baselines on the FisheyeNeRF Dataset [34] across three scenes. Vanilla 3DGS [37] fails completely to reconstruct the scenes because lens distortion is not accounted for. Fisheye-GS [46] adopts a parametric model, but the peripheral regions produce blurry results, as highlighted by the red box in the corners. In contrast, our method achieves clean and sharp reconstructions.
The second part of the video shows reconstruction results using our method of walk-around captures. These include both synthetic and real-world scenes. Our approach achieves sharp and clean renderings once completing the reconstruction.
The last part of the video compares visualization of our method and a conventional reconstruction pipeline that either uses narrow FOV perspective images or undistorted images from COLMAP [58] as input. Our method can successfully reconstruct larger regions, particularly for scenes captured with large 180 cameras. Besides, we render videos in fisheye views for each scene.
2 Synthetic Data Creation
To carefully control the experimental settings, we customized a camera module in the Mitsuba ray tracer [33] using camera parameters derived from DSLR lenses, as profiled in the open-source Lensfun [1]. We also utilize 3D assets, including geometry, materials, and lighting from [9] to produce renderings. Our synthetic dataset contains three large indoor scenes and four object-centric scenes. All indoor scenes follow a Sigma 180 circular fisheye camera. Two of the object scenes are rendered with a 120 fisheye lens, and the others are rendered with classic radial distortion.
To capture detailed perspectives of the scene, cameras are placed close to the objects at varying distances. For Room2, since objects were uniformly distributed in the space, we placed a set of camera centers along a Hilbert curve, then look out from each fix camera center to cover the surroundings. The number of images captured at each point is reduced by half for 180 images compared to those with 100 FOV.
In order to verify the accuracy of self-calibration, we also generate a set of hold-out cameras that share the same distribution as the training set. For each validation camera, we render paired perspective and 180 fisheye images. Unlike real-world datasets such as Garden and Studio, our synthetic dataset allows comparison with Ground-Truth perspective views. As shown in Fig. 8, we render an additional 20 for the hold-out cameras to show the difference in coverage between our method and conventional capture approach.
Room1 | ![]() |
Kitchen | ![]() |
Room2 | ![]() |
(a) Small FOV | (b) Ours | (b) GT |
3 Extra Comparisons with Baseline
In addition to Fig. 6 in the main paper, we also provide a quantitative evaluation of our method compared with the baseline Fisheye-GS [46]. The performance degradation observed in the baseline method is primarily due to the limitations of the parametric model used during reconstruction, which struggles with the edge of large FOVs. As a result, there is no geometric consistency in the peripheral regions to produce uniform gradients for optimizing the Gaussians. When rendering novel views, these Gaussians appear as large floaters that occlude the camera view. We further visualize this phenomenon in the failure case video reconstructed by Fisheye-GS [46]. The center area was revealed as we gradually decreased the scale of the Gaussians. Please refer to our supplementary videos under “section3_video” directory for the comparison with the parametric distortion model.
Method | Garden | Studio | ||||
SSIM | PSNR | LPIPS | SSIM | PSNR | LPIPS | |
Fisheye-GS [46] | 0.530 | 14.94 | 0.542 | 0.536 | 12.24 | 0.549 |
Ours | 0.882 | 27.85 | 0.144 | 0.965 | 33.86 | 0.044 |
(a) Low Resolution | (b) High Resolution |

### v2

@@ 3.3 Cubemap for Large FOV
In order to apply our method to cameras with larger FOV, we extend the single-planar perspective projection to a cubemap projection [60, 37]. Mathematically, single-planar projection requires upsampling in the peripheral regions, and the sampling rate increases drastically as the FOV approaches 180. In contrast, rendering with a cubemap maintains a relatively uniform pixel density from the image center to the edges, making it ideal for wide-angle rendering.
@@
Given the parameters estimated from SfM [59], the parametric model is then applied to reproject the raw image into perspective images, as shown in the blue box in Fig. 2(a). These reprojected images are then used for reconstruction through perspective-based pipelines like NeRF [50] or 3D Gaussian Splatting (3DGS) [38]. However, this process stretches the pixels in the peripheral regions, and the effect becomes more pronounced when the images are reprojected to larger FOV perspectives, as shown in the orange box. More specifically, the stretching rate of each pixel is defined by the inverse of Eq. 3, which exhibits a trend similar to , where is the FOV angle from the center of the raw image. When the FOV of the reprojected image is 110 (as in the blue example), the upsampling rate from the blue circle to the box is approximately 1.4. However, when the FOV increases to 170, as in the orange example, this rate increases to 11.4, inevitably sacrificing a significant amount of high-frequency information for reconstruction.
Moreover, to preserve central details, the resolution of undistorted images needs to be higher, as the pixel density at the center of the raw image should ideally match that of perspective ones. For example, when undistorting a fisheye image in Fig. 2(a), the resulting perspective image in the orange region would have an extremely high resolution, making it computationally expensive to render. A common solution is to crop away the periphery, following COLMAP’s solution [59], but this strategy contradicts our intention of using a fisheye camera to capture wide-angle information.
@@
Inspired by the representation of environment maps using cubemaps [25] and hemi-cube [17] in computer graphics, we propose representing extreme wide-angle renderings with cubemap projections, each covering 90 FOV and oriented orthogonally to one another, as illustrated in Fig. 2(b). By resampling across the cubemap faces, we can render perspective or distorted images with FOVs even larger than 180. Fast rasterization is first applied to obtain each face of the cubemap. For each rendered pixel, we look up its corresponding position in the constructed cubemap. Our hybrid distortion field then resamples from the lookup table to achieve the distorted rendering. In practice, it is not necessary to render all faces of the cubemap at once, as the number of cubemap faces may vary for different FOV camera lenses.
The resampling step involves only a simple coordinate transformation along with our hybrid field distortion. The distance from the shared camera center to each plane is normalized to 1. To render a pixel outside the 90 FOV at where and , for instance, the pixel on the right face can be obtained as in camera coordinates, looking toward the right side. When considering lens distortion, the sampling mapping is distorted according to Eq. 7, altering the lookup on the right face to , where . By doing so, the entire distortion field can be directly applied to the cubemap for large FOV rendering. The entire resampling process is fully differentiable, making it directly applicable in our hybrid distortion field as a plug-and-play module.
@@
3DGS [38] constructs an ordered set of Gaussians before alpha blending [38]. Gaussians are sorted based on their orthogonal projection distance to the image plane. This approach is valid as long as a single Gaussian is not projected onto multiple faces. However, with the cubemap representation, Gaussians can span the boundary between two faces, leading to multiple projections with inconsistent ordering across faces. This discrepancy introduces intensity discontinuities at the boundaries. To address this issue, we replace the original sorting strategy with a distance-based approach, ordering Gaussians by their distance from the camera center. This ensures that the rasterization order remains consistent across all cubemap faces, thereby alleviating intensity discontinuities. Due to the affine approximation used in [38, 68], the 2D covariance of Gaussians near cubemap face boundaries still has a slight influence on the final rendering, which we further discuss in Sec. 5 and the supplementary material.
3.4 Optimization of Camera Parameters
Our pipeline differentiates all camera parameters. We theoretically derive the gradient for all camera parameters, including extrinsics and intrinsics, making the camera module completely differentiable and capable of being optimized alongside distortion modeling. All gradient calculations are implemented with a native CUDA kernel. A comprehensive mathematical derivation and experimental results are provided in the supplementary.
@@ 3.4 Optimization of Camera Parameters
3DGS-perspective [38] | 0.431 | 14.06 | 0.547 | 0.507 | 15.21 | 0.533 | 0.281 | 12.91 | 0.609 | 0.502 | 15.09 | 0.530 | 0.505 | 15.19 | 0.549 | 0.297 | 12.70 | 0.595 |
3DGS-COLMAP [38] | 0.583 | 18.28 | 0.290 | 0.637 | 21.64 | 0.296 | 0.443 | 18.09 | 0.379 | 0.580 | 19.63 | 0.327 | 0.660 | 20.87 | 0.282 | 0.511 | 20.24 | 0.280 |
@@
4 Experiments
In this section, we first briefly introduce our data preparation pipeline in Sec. 4.1. We compare our method against various baselines for scene reconstruction with large-FOV input (Sec. 4.2 and Sec. 4.3), followed by ablation studies for both the neural distortion model and the cubemap projection in Sec. 4.4.
4.1 Data Acquisition
We customized a camera module in the Mitsuba ray tracer [35] to incorporate fisheye camera parameters derived from the open-source Lensfun database [1]. Using a 180 fisheye camera, we rendered three scenes from [9]. We generated a training set with both perspective and fisheye views for baselines and our method, respectively. Additionally, we captured several real-world datasets using different uncalibrated cameras to evaluate our method. These datasets consist of casual walk-around video footage captured using a camera with an approximate 150 FOV. Furthermore, we tested our method on the existing FisheyeNeRF benchmark dataset [36], which contains fisheye images with an FOV of approximately 120.
(a) Fisheye-GS | (b) Ours | (c) GT | (d) Fisheye-GS | (e) Ours | (f) GT |
@@ 4.1 Data Acquisition
4.2 Comparisons to Traditional Lens Models
We first validate that our hybrid field representation models large-FOV cameras better than traditional polynomial distortion models [10, 59]. Using the FisheyeNeRF dataset [36], we compare our method against four baselines. The first baseline is Vanilla 3DGS [38], which only uses a perspective camera model. Second, we directly apply the polynomial distortion model estimated by COLMAP [59] to 2D Gaussians after the rasterization of 3DGS [38]. Third, we evaluate Fisheye-GS [47], which modifies the projection of Gaussians to a specific fisheye parametric model. Finally, we re-implemented ADOP [57] using an omnidirectional camera model as the fourth baseline.
We report quantitative results in Tab. 1, and our method consistently outperforms all baselines. We also show novel-viewpoint renderings in Fig. 4. Baseline methods such as ADOP [57] and Fisheye-GS [47] tend to produce more artifacts in the corners of the test views, whereas our method reconstructs finer details.
To evaluate how our method performs with increasing FOV, we conducted an additional experiment on two datasets with different camera FOVs: real-world scenes captured with 150 cameras and a synthetic dataset with a 180 FOV. As the FOV approaches 180, as shown in Fig. 5, our method successfully handles peripheral regions, whereas Fisheye-GS [47] fails, producing spiky Gaussians in the surroundings and introducing artifacts at the center that blur the rendering. The quantitative evaluation of Fig. 5 can be found in the supplementary material.
@@ 4.2 Comparisons to Traditional Lens Models
Our method can also be applied to different types of lens distortion. We introduce radial and tangential distortions to images in the LLFF dataset [49], using camera parameters derived from the Lensfun database [1]. By applying these parameters, we generate a combination of distortions, as demonstrated in the T-Rex scene. We optimize our model using these distorted images and obtain the distortion map learned by our lens model. Fig. 6 shows that our method accurately recovers various types of camera distortion without manual calibration or access to the physical camera.
(a) 3DGS w/ COLMAP | (b) Ours | (c) 3DGS w/ Regular Camera | (d) Ours |
4.3 Large FOV Reconstruction with Few Captures
By enabling the use of large-FOV views, our method can reconstruct scenes with fewer images than would be needed for narrower perspective views. We evaluate our method on 150 real-world captures and 180 synthetic scenes. For real-world scenes, we follow the conventional paradigm of using COLMAP to first resample the images to perspective—including peripheral cropping performed by COLMAP [59]—before reconstruction. For synthetic scenes, we can flexibly modify the camera model and render images with a regular 90 FOV as input for 3DGS [38]. Since our method supports perspective rendering after training, we compare it against the baseline using a set of perspective hold-out cameras for a fair evaluation. Quantitative evaluation is conducted only on synthetic data, as obtaining ground truth perspective images for real-world captures is not feasible.
Method | Num | SSIM | PSNR | LPIPS |
3DGS [38] | 200 | 0.654 | 19.08 | 0.332 |
Ours | 100 | 0.800 | 29.01 | 0.231 |
50 | 0.735 | 26.26 | 0.267 | |
25 | 0.709 | 24.59 | 0.292 | |
10 | 0.615 | 22.77 | 0.356 |
We report quantitative results in Tab. 2, where our method outperforms the baseline even with far fewer input views, down to 10-15%. The trajectories of Room 1 and Garden are similar, both focusing inward, while Room 2 and Studio involve a walk-around motion covering the entire space. As a result, we observe large incomplete regions in the first row of Fig. 7 for the baseline, whereas our method achieves complete reconstruction. The walk-around scenes shown in the last row of Fig. 7 exhibit noticeable artifacts and floating elements, primarily due to small-FOV captures or cropping in COLMAP. These experiments demonstrate the efficiency of our approach compared to regular-FOV cameras, achieving high-quality reconstruction with significantly fewer captures, as shown in Tab. 2.
Explicit Grid | iResNet | PSNR | SSIM | LPIPS |
✗ | ✓ | Out-of-Memory | ||
✗ | ✗ | 14.19 | 0.421 | 0.561 |
✓ | ✗ | 19.79 | 0.569 | 0.309 |
✓ | ✓ | 23.67 | 0.777 | 0.151 |
Projection | PSNR | SSIM | LPIPS |
Single Plane | 24.10 | 0.676 | 0.312 |
Cubemap | 29.01 | 0.792 | 0.253 |
4.4 Ablation Studies
Hybrid Field.
To verify the effectiveness of each module in our hybrid field, we isolate individual components (either iResNet or the explicit grid) and evaluate their performance separately. We report quantitative results in Tab. 3 on FisheyeNeRF [36] scenes, showing that neither module alone achieves optimal performance compared to our full method. To test the grid alone, we use an explicit grid of learnable displacement vectors. This strategy provides only a slight improvement, primarily because the optimization of the explicit grid is prone to overfitting and can become trapped in local minima. We visualize the difference between the explicit grid and our hybrid field distortion flow in Fig. 3. We cannot report results for iResNet without the grid, as explained in the last paragraph of Invertible Residual Networks in Sec. 3.2. Our full method (iResNet + control grid) achieves the best performance by combining the expressiveness of iResNet with the computational efficiency of the explicit grid. This ablation verifies the necessity of our hybrid representation for modeling distortion.
Cubemap Resampling.
To validate the necessity of cubemap resampling for wide-FOV reconstruction, we compare results obtained using our hybrid distortion field with single-plane projection versus cubemap resampling. We evaluate the performance of these two methods in the Mitsuba synthetic scenes. As presented in Tab. 4, using cubemap projections produces significantly better results when reconstructing images captured by 180 fisheye lenses. Additional qualitative comparisons between single-plane and cubemap projections are provided in the supplementary.
@@ 5 Discussion
This paper presents a method for optimizing 3D Gaussian representations while self-calibrating camera parameters and lens distortion. Our approach enables the use of large field-of-view captures to achieve efficient and high-quality reconstruction without cumbersome pre-calibration. Even with fewer input captures, our method maintains comprehensive scene coverage and reconstruction quality.
@@
This work does not account for entrance pupil shift in fisheye lenses, which affects near-field scenes by causing splat misalignment but is negligible for distant scenes. Vignetting and chromatic aberration were also not modeled, leaving room for future extensions. Additionally, intensity discontinuities arise at cubemap face boundaries due to variations in 2D covariances. A simple fix involves switching from depth sorting to distance sorting, while a more refined solution could involve projecting covariances orthogonally to the viewing ray. More discussions can be found in the supplementary.
@@ 1 Supplementary Video
We provide a video, “supp_video.mp4,” to better compare our method with baselines. Our video is organized into three parts.
The first part presents a comparison between our method and baselines on the FisheyeNeRF dataset [36] across three scenes. Vanilla 3DGS [38] completely fails to reconstruct the scenes because lens distortion is not accounted for. Fisheye-GS [47] adopts a parametric model, but the peripheral regions produce blurry results, as highlighted by the red box in the corners. In contrast, our method achieves clean and sharp reconstructions.
The second part of the video shows reconstruction results using our method on walk-around captures, including both synthetic and real-world scenes. Our approach achieves sharp and clean renderings upon completing the reconstruction.
The third part of the video compares visualizations of our method with a conventional reconstruction pipeline that either uses narrow-FOV perspective images or undistorted images from COLMAP [59] as input. Our method successfully reconstructs larger regions, particularly for scenes captured with large 180 cameras. Besides, we render videos in fisheye views for each scene.
Additionally, we provide “opt_pose.mp4,” which illustrates the camera changes during optimization, and “fisheye-gs_failure.mp4,” which demonstrates failure cases of Fisheye-GS [47] in extremely large-FOV settings.
We strongly encourage all reviewers to watch the provided video for a more comprehensive visual understanding of our results.